9. Empire Electric Company (EEC) uses only debt and common equity. It can borrow unlimited amounts...

9.

Empire Electric Company (EEC) uses only debt and common equity. It can borrow unlimited amounts at an interest rate of rd = 10% as long as it finances at its target capital structure, which calls for 30% debt and 70% common equity. Its last dividend (D0) was $2.25, its expected constant growth rate is 4%, and its common stock sells for $29. EEC's tax rate is 25%. Two projects are available: Project A has a rate of return of 12%, and Project B's return is 10%. These two projects are equally risky and about as risky as the firm's existing assets.

  1. What is its cost of common equity? Do not round intermediate calculations. Round your answer to two decimal places.

      %

  2. What is the WACC? Do not round intermediate calculations. Round your answer to two decimal places.

      %

  3. Which projects should Empire accept?

    -Select-Project AProject B
Answer #1

a)

Year 1 dividend = 2.25 (1 + 4%) = 2.34

Cost of common equity = (D1 / price) + growth rate

Cost of common equity = (2.34 / 29) + 0.04

Cost of common equity = 0.08069 + 0.04

Cost of common equity = 0.1207 or 12.07%

b)

WACC = Weights * costs

WACC = 0.3*0.1*(1 - 0.25) + 0.7*0.1207

WACC = 0.0225 + 0.08449

WACC = 0.1070 or 10.70%

c)

Project A should be accepted as it has a higher return